# Mitigating Graph Covariate Shift via Score-based Out-of-distribution Augmentation

### v1

@@ 4 Score-based Out-of-distribution Graph Augmentation
$\nabla_{G_{t}}\log p_{t}\left(G_{t}\mid\mathbf{y}_{G},\mathbf{y}_{\text{ood }}=\lambda\right)=(1-\sqrt{\lambda})\nabla_{G_{t}}\log p_{t}\left(G_{t}\right)+(1-\sqrt{\lambda})\nabla_{G_{t}}\log p_{t}\left(\mathbf{y}_{G}\mid G_{t}\right)$ (9)
As seen in Equation 9, we need to compute the target class probability for the conditional score functions. To achieve this, we train a classifier $\phi_{t}$ to predict graph label $\mathbf{y}_{G}$ from the noisy graph $G_{t}$ at time step $t$: $\phi_{t}(G_{t})=\hat{\mathbf{y}_{G}}$. The output probability of $\phi_{t}$ can approximate the distribution $p_{t}\left(\mathbf{y}_{G}\mid G_{t}\right)$.
Consequently, the conditional score function in Equation 6 results in a marginal distribution proportional to $p_{t}\left(G_{t},\mathbf{y}_{G}\right)^{1-\sqrt{\lambda}}$. When $\lambda=0$, the marginal distribution in Equation 6 simplifies to $p_{t}\left(G_{t},\mathbf{y}_{G}\right)$. The reverse-time diffusion process denoises the perturbed graphs to the augmented distribution $P_{\mathrm{tr}}\left(G,Y\mid\mathbf{y}_{\text{ood }}=0\right)$, which closely resembles the original distribution $P_{\mathrm{tr}}\left(G,Y\right)$. As $\lambda$ increases, the augmented distribution $P_{\mathrm{tr}}\left(G,Y\mid\mathbf{y}_{\text{ood }}=\lambda\right)$ becomes more uniform relative to the original data distribution. By adjusting $\lambda$, we can flexibly control the dispersion, making the generated graphs more likely to be out-of-distribution. Consequently, we can increase the divergence between $P_{\mathrm{tr}}\left(G,Y\mid\mathbf{y}_{\text{ood }}=\lambda\right)$ and $P_{\mathrm{tr}}\left(G,Y\right)$.
Working principles of OODA. OODA not only explores controllable environmental features but also preserves stable features. According to Equation 7, the term $\nabla_{G_{t}}\log p_{t}\left(\mathbf{y}_{\text{ood }}=\lambda\mid G_{t},\mathbf{y}_{G}\right)$ directs the reverse diffusion process towards out-of-distribution regions with graphs that exhibit the explored environmental patterns. Additionally, the term $\nabla_{G_{t}}\log p_{t}\left(\mathbf{y}_{G}\mid G_{t}\right)$ guides the reverse diffusion process towards regions containing graphs that are highly likely to possess patterns determining the target class. Although we use a classifier to approximate the target class probability $p_{t}\left(\mathbf{y}_{G}\mid G_{t}\right)$, the classifier is able to predict the target class even in the presence of noise. By iteratively adding noise, the classifier captures the relationship between the stable patterns in the perturbed graph $G_{t}$ and the target class. Together, these two terms guide the sampling process to generate graphs that contain desired stable patterns and extend beyond the training environments.
In addition to the aforementioned principles, OODA can generate diverse and valid OOD graph instances for the final graph classification problem under distributional shifts. This capability is based on the inherent randomness in the forward processes and the effectiveness of the reverse process in the diffusion model.
Implementations of OODA. Directly applying this framework to graphs proved inadequate for capturing the intricate relationships between nodes and edges, essential for accurately learning graph distributions (Jo et al., 2022). To overcome this, we simultaneously models the diffusion processes of node features and adjacency matrices of perturbed graphs $\left\{G_{t}=\left(\bm{X}_{t},\bm{A}_{t}\right)\right\}_{t=0}^{T}$ using a set of SDEs for Equation 6:
[…]
where $\mathbf{f}_{t}(\bm{X},\bm{A})=\left(\mathbf{f}_{1,t}(\bm{X}),\mathbf{f}_{2,t}(\bm{A})\right)$ and $g_{t}=\left(g_{1,t},g_{2,t}\right)$ representing the drift and diffusion coefficients, respectively. The reverse-time processes are captured by standard Wiener processes $\overline{\mathbf{w}}_{1}$ and $\overline{\mathbf{w}}_{2}$, with $\mathrm{d}\bar{t}$ indicating an infinitesimally small negative time step. We train one graph transformer (Dwivedi & Bresson, 2020; Vignac et al., 2022), denoted as $s_{\theta,t}=\left(s_{\theta_{1},t},s_{\theta_{2},t}\right)$ to closely estimate the partial score functions $\nabla_{\bm{X}_{t}}\log p_{t}\left(\bm{X}_{t},\bm{A}_{t}\right)$ and $\nabla_{\bm{A}_{t}}\log p_{t}\left(\bm{X}_{t},\bm{A}_{t}\right)$, facilitating the backward simulation of the equation to simultaneously generate node features and adjacency matrices of unlabelled graphs. Therefore, the denoising graph transformer is only trained with unlabelled graphs without $\lambda$ values and graph labels. The graph transformer with perturbed node features $X_{t}$, adjacency matrices $A_{t}$ and normalized timestep as input. The timestep $t$ value is treated as a global graph feature, and an embedding layer is used to embed $t$.
We also use a graph transformer model $\phi_{t}$ with the same architecture to predict the class label of the noisy graphs $G_{t}=\left(\bm{X}_{t},\bm{A}_{t}\right)$ at time step $t$. The target class $j$ probability $p_{t}\left(\mathbf{y}_{G}=j\mid\bm{X}_{t},\bm{A}_{t}\right)$ is then given by:
$p_{t}\left(\mathbf{y}_{G}=j\mid\bm{X}_{t},\bm{A}_{t}\right)=\frac{e^{\phi_{t}\left(\bm{X}_{t},\bm{A}_{t}\right)_{[j]}}}{\sum_{j=1}^{M}e^{\phi_{t}\left(\bm{X}_{t},\bm{A}_{t}\right)_{[j]}}}$
Once both the score transformer and the classifier are trained, we use them to compute the conditional partial scores during the sampling process.
$\alpha_{1,t}$ $=0.1^{t}\frac{r_{1}\left\|\bm{s}_{\theta_{1},t}\left(\bm{G}_{t}\right)\right\|}{\left\|\nabla_{\bm{X}_{t}}\log p_{t}\left(\mathbf{y}_{G}\mid\bm{X}_{t},\bm{A}_{t}\right)\right\|}$ (11)
$\alpha_{2,t}$ $=0.1^{t}\frac{r_{2}\left\|\bm{s}_{\theta_{2},t}\left(\bm{G}_{t}\right)\right\|}{\left\|\nabla_{\bm{A}_{t}}\log p_{t}\left(\mathbf{y}_{G}\mid\bm{X}_{t},\bm{A}_{t}\right)\right\|}$
[…]
Intuitively, at the early stages of the reverse-time SDEs, the graphs are highly perturbed, resembling the prior noise distribution. Therefore, the classifier cannot accurately approximate the target class probability. Consequently, in the initial denoising steps, we focus more on guiding the reverse-time SDEs towards the OOD distribution. As the reverse-time SDEs progressively denoise the graphs, we introduce guidance to direct the reverse-time SDEs towards regions exhibiting the desired stable patterns and the explored OOD environmental patterns.
5 Experiments
In this section, we first demonstrate the effectiveness of our diffusion models on graph OOD tasks in Section 5.2, then validate the efficacy of our OOD-controlled diffusion process on GOOD-Motif and GOOD-HIV datasets in Section 5.3. We further conduct an ablation study to verify the effectiveness of our diffusion models to generate OOD graphs in Section 5.3
5.1 Experimental Settings
@@
Setup. For a fair comparison, we adopt the same evaluation metrics as those used in (Gui et al., 2022). The model that achieves the best performance on the OOD validation sets is then evaluated on the OOD test sets. Furthermore, to ensure fair comparison across all methods, we utilize the same GNN backbones—GIN (Xu et al., 2019) and GIN-Virtual (Xu et al., 2018; Gilmer et al., 2017)—as applied in the GOOD benchmark (Gui et al., 2022) for each dataset. The experimental details, including evaluation metrics and hyperparameter configurations, are summarized in Appendix A.4.
Datasets. We use synthetic, semi-artificial, and real-world datasets from GOOD (Gui et al., 2022), including GOOD-Motif, GOOD-CMNIST, GOOD-HIV, and GOOD-SST2. Consistent with (Gui et al., 2022), we apply base, size, color, scaffold, and length data splits to introduce diverse covariate shifts in graph structure, node features, and edge features. Detailed descriptions of the datasets are provided in Appendix A.4.
Baselines. We adopt 16 baselines, which can be divided into the following three specific categories:(i) general generalization algorithms, including ERM, IRM (Arjovsky et al., 2019), GroupDRO (Sagawa et al., 2019), VREx (Krueger et al., 2021), DANN (Ganin et al., 2016), Deep Coral (Sun & Saenko, 2016); (ii) graph generalization algorithms, including DIR (Wu et al., 2022), GSAT (Miao et al., 2022), CIGA (Chen et al., 2022), and (iii) graph data augmentation techniques, including DropNode (Feng et al., 2020), DropEdge (Rong et al., 2019), MaskFeature (Thakoor et al., 2021), FLAG (Kong et al., 2022), M-Mixup (Wang et al., 2021), G-Mixup (Han et al., 2022), AIA (Sui et al., 2024).
5.2 Graph Out-of-Distribution Classification
The graph classification performances under covariate shift are presented in Table 1. As shown, OODA consistently outperforms all baseline methods across diverse covariate shifts and different datasets.
On the synthetic dataset GOOD-Motif, OODA achieves a performance improvement of $6.59\%$ over ERM under base shift and $9.07\%$ under size shift. For the semi-artificial dataset GOOD-CMNIST, designed for node feature shifts, performance is significantly enhanced by $18.23\%$ compared to the leading graph augmentation method, AIA, and improved by $21.40\%$ over the best graph invariant learning method, DIR. In the real-world molecular dataset GOOD-HIV, where covariate shifts occur in graph structure, node features, and edge features simultaneously, OODA outperforms ERM by $2.09\%$ on scaffold shift and by $6.53\%$ on size shift. For the real-world natural language sentiment analysis dataset GOOD-SST2, while AIA is outperformed by MaskFeature by $0.31\%$, OODA exceeds MaskFeature by $0.69\%$.
These results demonstrate that no baseline graph invariant learning or graph data augmentation methods consistently outperform each other under various covariate shifts. OODA enhances environmental exploration by generating out-of-distribution (OOD) graphs in a controlled manner while preserving stable features. Consequently, OODA reliably improves performance across different datasets facing various covariate shifts.
@@ 5.2 Graph Out-of-Distribution Classification
OODA(Ours) $\mathbf{75.25\pm 3.84}$ $\mathbf{60.81\pm 7.80}$ $\mathbf{54.60\pm 2.27}$ $\mathbf{72.67\pm 1.28}$ $\mathbf{66.47\pm 2.29}$ $\mathbf{82.69\pm 0.28}$
Table 1: Performance on synthetic and real-world datasets. Bold numbers indicate the best performance, while the underlined numbers indicate the second best performance.
5.3 OOD controlled graph generation
@@
In this section, we present both qualitative and quantitative experiments to demonstrate the effectiveness of our out-of-distribution diffusion augmentation framework. The experiments are conducted using the GOOD-Motif-base and GOOD-HIV-scaffold datasets.
Figure 2: Distance between original and OOD graph distributions on GOOD-Motif.
Controlled OOD generation. We first validate that our framework can explore the space of the original distribution $P_{\mathrm{tr}}(G,Y)$ and generate an augmented distribution $\tilde{P}_{\mathrm{tr}}(G,Y)$ in a controlled manner. To evaluate the deviation of the augmented distribution from the training graphs, we employ existing random Graph Isomorphism Networks (GIN)-based metrics (Thompson et al., 2022), which are well-suited for graph generative models. These metrics are more expressive, and significantly reduce computational costs, particularly on large graph datasets. As recommended by (Thompson et al., 2022), we utilize the maximum mean discrepancy (MMD) with a radial basis function (RBF) kernel, which is widely recognized as a robust metric for measuring distributional differences. MMD RBF is used to measure the discrepancy between the augmented and training distributions. We vary $\lambda$ uniformly in the range $[0,1)$ with a step size of $0.1$ to generate ten augmented datasets, each containing the same number of graphs as the training set, and compute the MMD RBF between each augmented dataset and the original training dataset. A detailed description of the estimation procedure is provided in Appendix A.3. The results for GOOD-Motif with base covariate shift are presented in Figure 2. As shown in Figure 2, the MMD RBF between the original GOOD-Motif-base graph distribution $P_{\mathrm{tr}}(G,Y)$ and the augmented graph distribution $\tilde{P}_{\mathrm{tr}}(G,Y)$ increases as $\lambda$ grows. When $\lambda=0.0$, the MMD RBF is $0.072\pm 0.002$. As $\lambda$ increases to $0.9$, the MMD RBF becomes $1.5$ times larger than that at $\lambda=0.0$. These results demonstrate that OODA can generate OOD graph samples in a controlled manner by flexibly adjusting the value of $\lambda$.
Figure 3: Expected probabilities that the OOD GOOD-Motif graphs retain stable patterns.
Stable patterns preservation. we utilize a pretrained graph transformer $\phi_{t}$, which was trained on noisy graphs from the training distribution, to compute the probability of each augmented OOD graph sample possessing stable patterns that determine the target class, denoted as $p\left(\mathbf{y}_{G}\mid\tilde{G_{0}}\right)$, where $\mathbf{y}_{G}$ represents the target class of $\tilde{G_{0}}$. We then compute the expected value $\mathbb{E}_{\tilde{G_{0}}\sim\tilde{P}}\left[p\left(\mathbf{y}_{G}\mid\tilde{G_{0}}\right)\right]$ across the augmented distribution. This approach verifies that the explored graphs in $\tilde{P}_{\mathrm{tr}}(G,Y)$ preserve the stable patterns characteristic of the graphs in $P_{\mathrm{tr}}(G,Y)$. The results for GOOD-Motif with base covariate shifts are illustrated in Figure 3. As depicted, the expected probability $\mathbb{E}_{\tilde{G_{0}}\sim\tilde{P}}\left[p\left(\mathbf{y}_{G}\mid\tilde{G_{0}}\right)\right]$ consistently exceeds $0.95$ as $\lambda$ increases, demonstrating the capability of OODA to generate OOD graphs that retain stable patterns.
Visualization of OOD graphs. We further demonstrate the efficacy of OODA by visualizing the OOD graphs generated by our approach in Table 3. In this visualization, three label-determining motifs—house, cycle, and crane—are highlighted in red, while the three environmental base graphs in the training distribution—wheel, tree, and ladder—are indicated in green. As illustrated in Table 3, increasing $\lambda$ leads to gradual modifications in the structures of the base graphs, while the motifs remain preserved.
@@
Figure 4: (Left): Distance between the original GOOD-HIV graph distribution and the augmented graph distribution. (Right): The validity of the OOD molecules and the expected probabilities that the OOD GOOD-HIV molecules retain stable patterns.
OOD molecules generation. To further validate the effectiveness of OODA in generating valid OOD molecules, we assess both the exploration of new patterns and the preservation of stable features. The Fréchet ChemNet Distance (FCD) (Preuer et al., 2018) is employed to quantify the distance between the training and augmented distributions of molecules, based on the penultimate activations of ChemNet. Additionally, RDKit (Landrum et al., 2016) is used to evaluate the fraction of valid molecules. We also compute the $\mathbb{E}_{\tilde{G_{0}}\sim\tilde{P}}\left[p\left(\mathbf{y}_{G}\mid\tilde{G_{0}}\right)\right]$ to confirm that the OOD molecules retain stable patterns necessary for inhibiting HIV replication. The results, shown in Figure 4, demonstrate that as $\lambda$ increases, the FCD between the training molecules and the generated OOD molecules grows. Despite this, the OOD molecules consistently preserve stable patterns, with $\mathbb{E}_{\tilde{G_{0}}\sim\tilde{P}}\left[p\left(\mathbf{y}_{G}\mid\tilde{G_{0}}\right)\right]$ remaining above $0.97$, while maintaining a $100\%$ validity rate.
@@
Ablation study. In this section, we present experiments to evaluate the impact of exploration guidance ($\lambda$) and stable patterns preservation guidance ($\alpha$). The results are summarized in Table 2. As shown, incorporating $\alpha$ guidance improves the OOD performance of the diffusion model trained on unlabeled graphs by ensuring that the generated graphs retain the stable patterns that determine the labels. Without $\alpha$ guidance, a diffusion model guided only by $\lambda$ tends to push the augmented graphs into arbitrary out-of-distribution (OOD) regions, which negatively impacts performance to some extent. Ultimately, the combination of both $\alpha$ and $\lambda$ guidance enables the augmented distribution to capture both stable patterns and novel environmental patterns, resulting in the best overall performance.
@@
OODA $\mathbf{75.25\pm 3.84}$ $\mathbf{72.67\pm 1.28}$ $\mathbf{82.69\pm 0.28}$
Table 2: Performance of OODA w/o environmental exploration guidance and stable pattern preservation guidance on synthetic and real-world datasets. Bold numbers indicate the best performance.

### v2

@@ 4.4. Working Principles of CODA
$(1-\sqrt{\lambda})\Big(\underbrace{\nabla_{G_{t}}\log p_{t}(G_{t})}_{\text{label-free}}+\underbrace{\nabla_{G_{t}}\log p_{t}(\mathbf{y}_{G}\mid G_{t})}_{\text{$C$-only}}\Big).$
Projecting onto $S$-coordinates gives $\nabla_{S_{t}}\log p_{t}(\mathbf{y}_{G}\mid G_{t})=0$, so the label-guidance field has no component on $S$. Thus the $S$-dynamics are label-independent for all $\lambda$.
(ii) $\lambda=0$ anchors $C\!\to\!Y$. At $\lambda=0$, (12) reduces to standard class-conditional guidance $\nabla\log p_{t}(G_{t})+\nabla\log p_{t}(\mathbf{y}_{G}\mid G_{t})$ targeting $P_{\mathrm{tr}}(G\mid Y)$; consequently $\tilde{P}_{0}(\tilde{C}\mid\mathbf{y}_{G}=y)\approx P_{\mathrm{tr}}(C\mid y)$ and $\tilde{P}_{0}(Y\mid G)=\tilde{P}_{0}(Y\mid C)$.
(iii) Continuation in $\lambda$. For $0<\lambda\leq 1$, the overall field is scaled by $(1-\sqrt{\lambda})\geq 0$ and the label term remains $C$-only. Since the $S$-evolution is label-free, changing $\lambda$ alters how broadly we explore in $S$ but cannot introduce dependence of $Y$ on $S$ given $C$. Hence $P(Y\mid G)=P(Y\mid C)$ is preserved along the reverse-time flow for any $\lambda\in[0,1]$.
Therefore, for all $\lambda\in[0,1]$, $\tilde{P}_{\lambda}(Y\mid G)=\tilde{P}_{\lambda}(Y\mid C)$, i.e., the causal relation $C\to Y$ is preserved.
Proposition 4.2 (Spurious pattern reduction).
Let $\tilde{P}_{\lambda}$ be the generated joint induced by (12). For any fixed class $y$, write $\tilde{\pi}_{\lambda}(s\mid y):=\tilde{P}_{\lambda}(S=s\mid\mathbf{y}_{G}=y)$ and $\pi_{0}(s\mid y):=P_{\mathrm{tr}}(S=s\mid Y=y)$. Then for $0\leq\lambda_{1}<\lambda_{2}\leq 1$ and any measurable “top-mass” set $\mathcal{S}_{\tau}(y):=\{\,s:\pi_{0}(s\mid y)\geq\tau\,\}$, it holds that
$\tilde{\pi}_{\lambda_{2}}\big(\mathcal{S}_{\tau}(y)\mid y\big)\;\leq\;\tilde{\pi}_{\lambda_{1}}\big(\mathcal{S}_{\tau}(y)\mid y\big).$
In words: as $\lambda$ increases, the probability mass assigned to training-typical spurious patterns $S$ monotonically decreases under $\tilde{P}_{\lambda}$.
Proof. (i) Semantics are $C$-anchored. By Section 4.1, $S\perp\!\!\!\perp Y\mid C$, hence $P(Y\mid G)=P(Y\mid C)$. Moreover the time-dependent classifier is trained on noisy pairs so that $p_{t}(\mathbf{y}_{G}\mid G_{t})=p_{t}(\mathbf{y}_{G}\mid C_{t})$, implying $\nabla_{S_{t}}\log p_{t}(\mathbf{y}_{G}\mid G_{t})=0$. Thus the label-guidance term in (12) has no component on $S_{t}$.
(ii) Low-density reweighting induces power tempering. With the OOD surrogate $p_{t}(\mathcal{E}_{\mathrm{ood}}=\lambda\mid G_{t},\mathbf{y}_{G})\propto p_{t}(G_{t},\mathbf{y}_{G})^{-\sqrt{\lambda}}$, the class-conditional sampler satisfies
$p_{t}(G_{t}\mid\mathbf{y}_{G},\mathcal{E}_{\mathrm{ood}}=\lambda)\;\propto\;p_{t}(G_{t}\mid\mathbf{y}_{G})^{\alpha(\lambda)},$
$\alpha(\lambda):=1-\sqrt{\lambda}\in(0,1].$
Marginalizing $C_{t}$ gives, for any fixed $y$,
$\tilde{\pi}_{\lambda}(s\mid y)\;\propto\;\int p_{t}(C_{t},s\mid y)^{\alpha(\lambda)}\,\mathrm{d}C_{t},$
followed by normalization over $s$. Note that $\alpha(\lambda)$ decreases as $\lambda$ increases.
(iii) Power tempering compresses high-probability advantages. Let $q_{s}:=\int p_{t}(C_{t},s\mid y)\,\mathrm{d}C_{t}=\pi_{0}(s\mid y)$. For any $s_{1},s_{2}$ with $q_{s_{1}}\geq q_{s_{2}}>0$ and any $\alpha\in(0,1]$,
$\frac{q_{s_{1}}^{\alpha}}{q_{s_{2}}^{\alpha}}=\Big(\frac{q_{s_{1}}}{q_{s_{2}}}\Big)^{\alpha}\leq\frac{q_{s_{1}}}{q_{s_{2}}}.$
Hence, as $\alpha(\lambda)$ decreases (i.e., $\lambda$ increases), the normalized mass shifts away from high-density (“typical”) $S$ configurations toward lower-density ones.
(iv) Conclusion. Combining (i)–(iii), increasing $\lambda$ cannot alter the $C$-anchored semantics, but it reduces the normalized probability of training-typical $S$ under $\tilde{\pi}_{\lambda}(\cdot\mid y)$, yielding the stated monotonicity on any top-mass set $\mathcal{S}_{\tau}(y)$.
5. Experiments
We first demonstrate the effectiveness of our diffusion models on graph OOD tasks. Then, we validate that the desired augmentation principles are successfully realized. Additional experiments regarding sensitivity of hyperparameter $\lambda$, as well as time and memory complexity, are presented in Appendix F and H.
5.1. Experimental Settings
@@
Setup. We adopt the same evaluation metrics as in (gui2022good) for a fair comparison. The model that achieves the best performance on the OOD validation sets is then evaluated on the OOD test sets. Furthermore, to ensure fair comparison across all methods, we utilize the same GNN backbones—GIN (xu2018how) and GIN-Virtual (xu2018powerful; gilmer2017neural)—as applied in the GOOD benchmark (gui2022good) for each dataset. The experimental details, including evaluation metrics and hyperparameter configurations, are summarized in Appendix B and D.
Table 1. Performance on synthetic and real-world datasets. Bold numbers indicate the best performance, while the underlined numbers indicate the second best performance. Bootstrap test is used to assess significance with $20000$ resamples. The dagger symbol $\left({}^{\dagger}\right)$ indicates that our method significantly outperforms the best baseline with a p-value less than $0.05$.
@@
CODA(Ours) $\mathbf{75.25\pm 3.84}^{\dagger}$ $\mathbf{60.81\pm 7.80}^{\dagger}$ $\mathbf{54.60\pm 2.27}^{\dagger}$ $\mathbf{72.67\pm 1.28}^{\dagger}$ $\mathbf{63.66\pm 1.20}^{\dagger}$ $\mathbf{82.69\pm 0.28}^{\dagger}$
Datasets. We evaluate on synthetic, semi-artificial, and real-world benchmarks from GOOD (gui2022good): GOOD-Motif, GOOD-CMNIST, GOOD-HIV, and GOOD-SST2. Following the official protocol, we use the base, size, color, scaffold, and length splits to induce diverse covariate shifts in structure, node features, and edge features. Although some datasets are not Web-native, the shifts emulate common Web scenarios: motif and size shifts mirror evolving community wiring and graph growth in social/hyperlink networks; color shifts stand in for feature/style changes (e.g., UI or content modality drift); scaffold splits mimic cold-start categories/items in e-commerce or knowledge graphs; and length shifts capture post/review length drift in social and content platforms. These settings therefore serve as Web-analogous stress tests for OOD generalization on graphs. Full dataset statistics, preprocessing details, and split constructions are provided in the Appendix D.
Baselines. We adopt 17 baselines, which can be divided into the following three specific categories:(i) general generalization algorithms, including ERM, IRM (arjovsky2019invariant), GroupDRO (sagawa2019distributionally), VREx (krueger2021out), DANN (ganin2016domain), Deep Coral (sun2016deep); (ii) graph generalization algorithms, including DIR (wu2022discovering), GSAT (miao2022interpretable), CIGA (chen2022learning), and (iii) graph data augmentation techniques, including DropNode (feng2020graph), DropEdge (rong2019dropedge), MaskFeature (thakoor2021large), FLAG (kong2022robust), M-Mixup (wang2021mixup), G-Mixup (han2022g), AIA (sui2024unleashing), GRATIN (abbahaddou2025graph).
5.2. Graph Out-of-Distribution Classification
The graph classification performances under covariate shift are presented in Table 1. As shown, CODA consistently outperforms all baseline methods across diverse covariate shifts and different datasets. On the synthetic dataset GOOD-Motif, CODA achieves a performance improvement of $6.59\%$ over ERM under base shift and $9.07\%$ under size shift. For the semi-artificial dataset GOOD-CMNIST, designed for node feature shifts, performance is significantly enhanced by $18.23\%$ compared to the leading graph augmentation method, AIA, and improved by $21.40\%$ over the best graph invariant learning method, DIR. In the real-world molecular dataset GOOD-HIV, where covariate shifts occur in graph structure, node features, and edge features simultaneously, CODA outperforms ERM by $3.09\%$ on scaffold shift and by $3.72\%$ on size shift. For the real-world sentiment analysis dataset GOOD-SST2, while AIA is outperformed by MaskFeature by $0.31\%$, CODA exceeds MaskFeature by $0.69\%$.
We assess statistical significance using a paired nonparametric bootstrap test (davison1997bootstrap) at the $5\%$ level to compare CODA with leading baselines; details are in Appendix G. Overall, these results demonstrate that no baseline graph invariant learning or graph data augmentation methods consistently outperform each other under various covariate shifts. CODA enhances environmental exploration by generating OOD graphs in a controlled manner while preserving stable features. Consequently, CODA reliably improves performance across different datasets facing various covariate shifts.
5.3. OOD Controlled Graph Generation
@@
In this section, we present both qualitative and quantitative experiments to demonstrate the effectiveness of our OOD diffusion augmentation framework. The experiments are conducted using the GOOD-Motif-base, GOOD-SST2-length and GOOD-HIV-scaffold datasets. Additionally, we present a comparative analysis of CODA against various augmentation methods, evaluating their respective capacities to explore environments while preserving label-determining stable structures. We further provides visualizations of the OOD graphs generated by CODA for the GOOD-Motif-basis and GOOD-Molhiv-scaffold datasets.
Training Distribution Augmentation with Control. We first validate that our framework can explore the space of the original distribution $P_{\mathrm{tr}}(G,Y)$ and generate an augmented distribution $\tilde{P}_{\mathrm{tr}}(G,Y)$ in a controlled manner. To quantify deviation from training graphs, we adopt the GIN-based metrics of (thompson2022on), which are expressive and computationally efficient for graph generators. Following their recommendation, we report Maximum Mean Discrepancy (MMD) with an RBF kernel as a robust measure of distributional shift and diversity. We vary $\lambda\in[0,1)$ in steps of $0.1$ to generate ten augmented sets (each the same size as the training set) and compute MMD (RBF) between each augmented set and the original training set. Estimation details are in the Appendix.
GOOD-Motif-basis. Figure 2(a) shows that MMD (RBF) between $P_{\mathrm{tr}}(G,Y)$ and $\tilde{P}_{\mathrm{tr}}(G,Y)$ increases monotonically with $\lambda$. At $\lambda=0.0$, MMD is $0.072\pm 0.002$; by $\lambda=0.9$, it is approximately $1.5\times$ larger. Thus, CODA generates OOD graphs in a tunable manner by adjusting $\lambda$, while also increasing diversity as the exploration strength grows.
GOOD-SST2-length. Here, $\lambda$ controls OOD size (e.g., $\lambda=0.1$ adds one unit to graph size relative to training). We again sweep $\lambda$ in $[0,1)$ by $0.1$ to obtain ten augmented sets. As shown in Figure 2(b), MMD (RBF) rises with $\lambda$: from $0.032\pm 0.005$ at $\lambda=0.0$ to roughly $25\times$ that value at $\lambda=0.9$.
Stable Pattern Preservation. We use a pretrained graph transformer $\phi_{t}$, trained on noisy graphs from $P_{\mathrm{tr}}$, as a proxy to assess whether augmented graphs retain label-determining stable patterns. Specifically, we compute $p(\mathbf{y}_{G}\mid\tilde{G}_{0})$ and report $\mathbb{E}_{\tilde{G}_{0}\sim\tilde{P}}[p(\mathbf{y}_{G}\mid\tilde{G}_{0})]$. On GOOD-Motif-basis (Figure 2(c)), this expectation remains above $0.95$ across $\lambda$; on GOOD-SST2-length (Figure 2(d)), it exceeds $0.94$. These results indicate that CODA generates OOD graphs while preserving stable, label-determining structure.
Comparison with augmentation baselines. Figure 2 compares methods on (i) environmental exploration and (ii) preservation of stable structures. G-Mixup mixes continuous labels, making discrete structure preservation ill-posed; we therefore omit that metric for G-Mixup. Chemistry-specific metrics (e.g., FCD (preuer2018frechet), RDKit-based scores (landrum2016rdkit)) require valid discrete molecules and are incompatible with baselines that output continuous graph features (e.g., AIA), so we focus on general graph metrics. Across GOOD-Motif-basis and GOOD-SST2, DropEdge explores more aggressively than G-Mixup and AIA but often distorts stable patterns. In contrast, CODA achieves stronger, tunable exploration via $\lambda$ (Figure 2(a) and Figure 2(b)) while consistently preserving stable structures better than the baselines (Figure 2(c) and Figure 2(d)).
Figure 2. Augmentation comparison on GOOD-Motif-basis and GOOD-SST2-length. (a) GOOD-Motif-basis: distribution distance between $P_{\mathrm{tr}}(G,Y)$ and $\tilde{P}_{\mathrm{tr}}(G,Y)$ (MMD-RBF). (b) GOOD-SST2-length: distribution distance (MMD-RBF). (c) GOOD-Motif-basis: stable pattern preservation probability (probability augmented graphs preserve label-determining stable patterns). (d) GOOD-SST2-length: stable pattern preservation probability.
@@
Figure 3. (Left): Distance between the original GOOD-HIV graph distribution and the augmented graph distribution. (Right): The validity of the OOD molecules and the expected probabilities that the OOD GOOD-HIV molecules retain stable patterns.
[…]
@@
Ablation study. We present experiments to evaluate the impact of exploration guidance ($\lambda$) and stable patterns preservation guidance ($\alpha$). The results are summarized in Table 2. As shown, incorporating $\alpha$ guidance improves the OOD performance of the diffusion model trained on unlabeled graphs by ensuring that the generated graphs retain the stable patterns that determine the labels. Without $\alpha$ guidance, a diffusion model guided only by $\lambda$ tends to push the augmented graphs into arbitrary OOD regions, which negatively impacts performance to some extent. Ultimately, the combination of both $\alpha$ and $\lambda$ guidance enables the augmented distribution to capture both stable patterns and novel environmental patterns, resulting in the best overall performance.
Table 2. Performance of CODA w/o environmental exploration guidance and stable pattern preservation guidance on synthetic and real-world datasets. Bold numbers indicate the best performance.
@@
CODA $\mathbf{75.25\pm 3.84}$ $\mathbf{72.67\pm 1.28}$ $\mathbf{82.69\pm 0.28}$
Visualization of OOD graphs We further demonstrate the efficacy of CODA by visualizing the OOD graphs generated by our approach in Table 3. In this visualization, three label-determining motifs—house, cycle, and crane—are highlighted in red, while the three environmental base graphs in the training distribution—wheel, tree, and ladder—are indicated in green. As illustrated in Table 3, increasing $\lambda$ leads to gradual modifications in the structures of the base graphs, while the motifs remain preserved.
Table 3. Visualizations of the augmented GOOD-Motif-base graphs generated by CODA. The graphs $\tilde{G}$ with $\lambda=0.1$, $\lambda=0.2$, and $\lambda=0.3$ represent the OOD graphs generated by CODA under different values of $\lambda$.